A simple pendulum with solid metal bob has a time period T  . What will be the period of the same pendulum if it is made to oscillate in a non-viscous liquid of density one-tenth of the metal of the bob?

Grade:11

1 Answers

vikas askiitian expert
13 years ago

T = 2pi(L/g)1/2        ...........1

finally weight of body is reduced in water = (dbVbg - dLVbg)

db,dL are dencity is bob & liquid , Vb = volume of bob

mgeff = Vbg(db-dL)

dL = db/10          (given)

so , mgeff  = Vbdbg(9/10)

       geff = 9g/10  

time period = T1 = 2pi(L/geff)1/2           ................2

divide 2 by 1

T1/T = (10/9)1/2

T1 = (10/9)1/2T

approve my ans if u like it